### v1

@@ Abstract
The European legislature has proposed the Digital Services Act (DSA) and Artificial Intelligence Act (AIA) to regulate platforms and Artificial Intelligence (AI) products. We review to what extent third-party audits are part of both laws and to what extent access to models and data is provided. By considering the value of third-party audits and third-party data access in an audit ecosystem, we identify a regulatory gap in that the Artificial Intelligence Act does not provide access to data for researchers and civil society. Our contributions to the literature include: (1) Defining an AI audit ecosystem that incorporates compliance and oversight. (2) Highlighting a regulatory gap within the DSA and AIA regulatory framework, preventing the establishment of an AI audit ecosystem. (3) Emphasizing that third-party audits by research and civil society must be part of that ecosystem and demand that the AIA include data and model access for certain AI products. We call for the DSA to provide NGOs and investigative journalists with data access to platforms by delegated acts and for adaptions and amendments of the AIA to provide third-party audits and data and model access at least for high-risk systems to close the regulatory gap. Regulations modeled after European Union AI regulations should enable data access and third-party audits, fostering an AI audit ecosystem that promotes compliance and oversight mechanisms.
@@
Auditing, Accountability, Data Access, AI Act, DSA, Regulation
@@ 1 Introduction
Artificial Intelligence (AI) technologies can cause serious harms to individuals, groups, and society. These technologies and products include Machine Learning (ML)-based algorithmic decision-making (ADM) systems, social media recommendation systems and general-purpose AI including foundational and generative AI. For instance, discriminatory practices and opaqueness in the welfare fraud ADM of the city of Rotterdam were uncovered [17]. It was demonstrated that YouTube [74] progresses towards extreme right-winged content on the platform. Furthermore, we have seen harmful biases in ML models such as stable diffusion [59], face recognition [13] or ChatGPT [72]. The European legislature has responded with different regulatory approaches to address these risks. For example, it has proposed regulating very large online platforms [^1] Not restricted to AI-driven algorithms (VLOPs) and very large online search engines (VLOSEs) via the Digital Services Act (DSA), as well as AI-driven products, services, and systems in the territory of the European Union (EU) via the Artificial Intelligence Act (AIA).
Among other regulatory tools, internal or external systematic evaluations of an algorithm can create public accountability by uncovering problematic machine behavior [70, 8, 68] and are to different extends part of the DSA and AIA. It has been argued that the EU regulation thereby de facto sketches an EU-wide ecosystem for auditing AI systems (see e.g. [56]). However, others have argued that, as enforcement mechanisms for “high-risk” AI systems, self-assessments before deployment (first- and second-party ex ante audits) and post-market monitoring (first- and second-party ex post audits) do not provide sufficient oversight [21]. Therefore, some have raised the question of the extent to which audits and impact evaluations by third parties (regulator, research, and civil society) are necessary for sufficient oversight [21, 44]. As this paper will show, access to the model and training data of the algorithmic system is crucial to facilitate third-party auditing. In this respect, AlgorithmWatch’s concern that this third party access is lacking in the AIA[4] should be taken seriously. The DSA includes third-party access by vetted researchers. However, others have argued that companies could leverage their market power against third-party auditors (i.e. audit capture) [46].
@@
In general, researchers wonder “who will audit internal auditors?” [19, 3, 21] Therefore, we analyze whether the DSA and the AIA can create an EU audit ecosystem. To do so, we examine the importance of third-party audits by researchers and civil society. We find that third-party audits are essential for the establishment of an AI auditing ecosystem in the EU that ensures accountability by compliance and oversight. We argue that the inclusion of such audits must be strengthened by considering the existing provisions of the DSA and the proposed provisions of the AIA. The question of how regulatory inspection should be included unfortunately goes beyond the scope of the paper, although researchers have justifiably insisted on the importance of this issue (see e.g. [36]).
@@
We contribute to the existing literature concerning third-party audits in three ways: (1) We define an AI audit ecosystem that accounts for compliance and oversight. (2) We demonstrate the existence of a regulatory gap as the DSA and AIA will not lead to an AI audit ecosystem. (3) We emphasize that third-party audits by research and civil society have to be part of that ecosystem and we demand that the AIA should include data and model access for certain AI products.
@@
First, we introduce the terms related to algorithmic audits and the different types of audits (section 2). We then analyze the current regulatory framework on third-party audits (section 3). The following section addresses third-party audit case studies and defines an AI audit ecosystem that ensures compliance and oversight. Furthermore, we show that third-party audits by researchers and civil society are vital for an AI audit ecosystem (section 4). Finally, we show that the current EU legislation – the DSA and AIA in combination – will not be sufficient to create an AI audit ecosystem (section 5).
@@ 2 Algorithm Audits: Terminology and Key Properties
In this section, our primary objective is to establish a robust foundation of terminology and key attributes inherent to audits. Consequently, we will distinguish audit types, examine the scope of audit types, and underline the role of sociotechnical system thinking in audits. These examinations are conducted in preparation for the analysis of the role of third parties in the DSA and AIA. Additionally, we will revisit these audit types in the discussion of creating a diverse AI audit ecosystem.
@@
The term “audit”  [^2] Sandvig et al. [76] have noted that ”although the word ’audit’ may evoke financial accounting, the original audit studies were developed by government economists to detect racial discrimination in housing.” Later, they were introduced in the social science as field experiments to test for discriminatory behavior [29] is used in various ways. For algorithmic audits [^3] Audit and algorithm audit will be used interchangeably in the following, some (e.g. Koshiyama et al. [44]) have provided a more technical definition of an algorithm audit, which is related to verification and compliance, and others (e.g. Sandvig et al. [76] and Metaxa et al. [53]) have used the term to describe a narrowly but still systematically targeted test of a particular aspect of a system (e.g., bias), originating from social science “audit studies.”
@@
Two examples demonstrate the various ways that the term has been used. The first, Bandy [8] identifies an audit research gap regarding Twitter audits. Twitter is the most researched social media platform due to its formerly openly-accessible API [62]. This suggests a conceptual problem: What distinguishes research on Twitter and an from an audit conducted on Twitter? In the second example Metaxa et al. [53] have defined audits that occur through systematic external evaluations by third parties as a means of activism with internal knowledge of the process or system being studied.
@@
We, however, will use audits as an umbrella term for empirical studies that investigate algorithmic systems and AI systems in particular and evaluate performance relative to expected behavior, as part of a broader accountability process [70]. Bandy [8] captures this broad definition, defining an algorithmic audit as “empirical study investigating a public algorithmic system for potential problematic behaviour” that includes the mentioned notions and is agnostic to the time of the audit (before or after the deployment). “Potential problematic behavior” refers to behavior with the potential to cause harm and includes the potential harms described by a sociotechnical AI system. ”Empirical studies” refers to qualitative and quantitative, as system expectations can be articulated and assessed in either mode [70].
@@
Bandy [8], who performed a systematic review of algorithm audits, notes that racial discrimination remains central to algorithm audits. Although so-called bias audits can account for algorithmic discrimination based on other protected attributes, such as age, gender, socioeconomic status, religion, and intersectional identities. In addition to such bias audits that can account for representational and allocative harms, other harm potentials are possible from AI systems. Therefore, audits might have different scopes to account for these harms, which include quality of service harms, interpersonal harms, and social system harms (e.g. Shelby et al. [79]). To account for these harms, audits like robustness audits, interpretability and explainability audits, security audits and privacy audits and can be conducted [44].
@@
  • first-party audits: internal audits which are conducted in an organization that developed the AI system based on specific metrics, toolkits, and requirements.
  • second-party audits: conducted by contractors with the developer (audits-as-a-service), including consulting companies as the “big four,” as well as companies that specialize in audits
@@
First-and second-party assessments and audits Third-party assessments and audits
Algorithmic risk assessment, Algorithmic impact assessment
Ethic-based auditing
Conformity assessment
Research and civil society algorithm audits, bias audits, adversarial audits
Regulatory inspection and certification
Algorithmic impact evaluations, sociotechnical audits
Example(s)
Case study algorithmic impact assessment for data access in a healthcare context [33], data protection impact assessment [11]
End-to-end framework for internal algorithmic auditing by Raji et al. [69]
AIA conformity assessments, example procedure by [28]; Medical Device Regulation Application Procedure [16]
Gender Shades [13], Audit of the Rotterdams’ welfare fraud ADM [17, 18], Audit Austrian AMS algorithm [5]
UK Information Commissioner’s Office’s ‘Guidance on the AI auditing framework’[41]
Stanford’s Impact evaluation of a predictive risk modeling tool [31], sociotechnical audit that assess police use of facial recognition[66]
Auditors
Creators or commissioners of the algorithmic system, contractors such as consulting companies
Creators or commissioners of the algorithmic system, contractors such as consulting companies
Creators or commissioners of the algorithmic system, contractors such as consulting companies
Researchers, investigative journalists, data scientists, NGOs, affected communities, or other stakeholders
Regulator or other auditing and compliance professionals that have an institutional role
Researchers, policymakers, investigative journalists, data scientists, NGOs, affected communities, or other stakeholders
Target Identification & Audit Scope
Broad scope: focus on the risks and negative consequences of the deployment of an algorithmic system
Narrow scope: focus predominantly on technical prevention and uncovering of specific harms (e.g. bias audit)
Broad scope: focus on an algorithmic system’s compliance with regulation and mostly non-technical and process-oriented
Narrow scope: focus predominantly on technical prevention and uncovering of specific harms (e.g. bias audit)
Broad scope: focus on an algorithmic system’s compliance with regulation
Broad scope: assessing possible societal impacts of an algorithmic system on the users or population it affects after it is in use
Auditor Independence
Internal and thus no independence or contractual dependence
Internal and thus no independence or contractual dependence
Internal and thus no independence or contractual dependence
External by independent organizations with no obligation
@@
External by independent institution, non-genuine independence when auditing administrative state-run ADMs
External by independent organizations with no obligation
Auditor Professionalization
Formal: done by deployers, internal teams or accredited second-parties
Formal: Done by deployers, internal teams or accredited second-parties
Formal: done by deployers, internal teams or accredited second-parties
@@
Mostly informal: accreditation and cooperation with company possible , although basic access could be possible without accreditation (’right to scrape’)
Formal: Accreditation
Mostly informal: accreditation and cooperation with company possible (vetted researchers in DSA)
When to audit
Before deployment
Before deployment
@@
Before deployment, post-market monitoring is possible
After deployment
After or before deployment, potentially an ongoing process
After deployment
@@
Table 1: This table is inspired by and derived from [2, 53, 57, 70] and should give an overview of assessments and audits that we surveyed. The considerations of Raji et al. [70] were used to display the differences of these assessment and audit types. However, this table claims neither completeness nor complete adequacy, as this topic is a relatively emerging one with respect to algorithmic systems. At the same time, a systematic presentation of assessments and audits differences is not the focus of this paper, even though we conceptually argue and advocate a broad deployment of assessments and audits in the AI lifecycle with the AI audit ecosystem.
@@
Algorithms cannot be divorced from the contexts in which they are applied [90] and represent a complex network containing economic, environmental, political, and social interests and outcomes affecting their functioning. The entanglement of technical components, and social elements leads to the typical acknowledgment of AI systems as sociotechnical systems [^4] Due to a broader understanding of their ecological impacts in recent years, [71] has added an “ecological” element to this notion, meaning that AI systems are socioecological-technical systems. Thus, it may be necessary to include ecological harms that AI causes in audits. [77] And, despite the common perception of audits as predominantly technical (see e.g. Koshiyama et al. [44]), internally conducted impact assessment as well as externally conducted impact evaluations and sociotechnical audits encompass a broader spectrum. They evaluate the impact of an algorithmic system on a population in a comprehensive framework that encompasses technical elements – most of the time limited technical elements – and include the social elements of the system [2]. We will come back to the assessment of sociotechnical systems in section 4.4.1.
@@
Ethics-based internal auditing and adversarial audits by researchers and civil society have a more technical scope and tend to focus on specific harms, e.g. bias audits. Nevertheless, drawing a clear line of demarcation between narrow and broad scope of audit is difficult. While maintaining a however central focus on sociotechnical systems, it is essential to consider the data that substantiate claims concerning specific harms attributed to an AI system. An example of this difficult distinction is that Stahl et al. [82] includes bias mitigation assessments in the systematic review on impact evaluations, which we have introduced as ethics-based audits or bias audits.
@@
Conformity assessments [56] and regulatory inspection [2] are systematic tests that evaluate if a system is compliant with a specification, standards, or a regulation. Regulatory inspections - conducted by an independent state-run body - are external to the deploying organization. Conformity assessments can be internal or external. In contrast to third-party audits, conformity assessments aim to achieve compliance, while third-party audits by researchers and civil society aim to prevent potential harm or uncover it by systematically evaluating the systems and checking the documents.
@@
We focus on third-party audits, especially those by researchers, academics, and journalists, as well as broader sociotechnical audits. Such audits are characterized by being conducted post-deployment through systematic evaluations, typically with little access to and independent of the organization deploying the algorithm [53, 19]. Thus, they have no formal contractual relationship with the audit target [67].

### v3

@@ Abstract
The European legislature has proposed the Digital Services Act (DSA) and Artificial Intelligence Act (AIA) to regulate platforms and Artificial Intelligence (AI) products. We review to what extent third-party audits are part of both laws and how is access to information on models and the data provided. By considering the value of third-party audits and third-party data access in an audit ecosystem, we identify a regulatory gap in that the AIA does not provide access to data for researchers and civil society. Our contributions to the literature include: (1) Defining an AI audit ecosystem incorporating compliance and oversight. (2) Highlighting a regulatory gap within the DSA and AIA regulatory framework, preventing the establishment of an AI audit ecosystem that has effective oversight by civil society and academia. (3) Emphasizing that third-party audits by research and civil society must be part of that ecosystem, we call for AIA amendments and delegated acts to include data and model access for certain AI products. Furthermore, we call for the DSA to provide NGOs and investigative journalists with data access to platforms by delegated acts and for adaptions and amendments of the AIA to provide third-party audits and data and model access, at least for high-risk systems, to at least reduce the regulatory gap that exists in the EU. Regulations modeled after the bloc’s AI regulations should enable data access and third-party audits, fostering an AI audit ecosystem that promotes compliance and oversight mechanisms.
@@
Auditing, Accountability, Data Access, AI Act, DSA, Black-box Audits, AI
@@ 1 Introduction
Artificial intelligence (AI) technologies can be found in technologies and products such as machine learning (ML)-based automated decision-making (ADM) systems, such as social media recommendation systems, and general-purpose AI, which includes generative AI models. These technologies have been proven to cause severe harm to individuals, groups, and society, especially through their capacity to enable discriminatory practices and for their opaqueness, such as in Rotterdam’s welfare fraud, whereby ADM systems deployed were biased against immigrants, who were frequently labeled as fraudsters against the system [20]. Scholars have also argued that YouTube’s algorithmic-based recommender systems [80] were more prone to recommend extreme content on the platform, helping spread disinformation and harmful content. Furthermore, harmful biases have been uncovered in ML models such as in the generative AI Stable Diffusion [65], face recognition APIs [14], and ChatGPT [78].
The European legislature has responded with different regulatory approaches to address these risks. For example, it has proposed regulating very large online platforms  [^1] Not restricted to AI-driven algorithms (VLOPs) and very large online search engines (VLOSEs) via the Digital Services Act (DSA) [31], as well as AI-driven products, services, and systems in the territory of the European Union (EU) via the Artificial Intelligence Act (AIA) [30].
Among other regulatory tools, both the DSA and the AI Act have provisions related to systematic evaluations of AI systems – denoted by the term audit – by agents that are both internal or external to the organisation developing or deploying these technologies and which aim at creating public accountability regarding these machine’s behavior [76, 8, 74]. These provisions include self-assessments both before deployment of the system — as is the case in the AI Act for “high-risk” AI systems which have to undergo a conformity assessment prior to be place in the EU Market [24] — and after deployment — as are the annual independent audits that VLOPs and VLOSEs have to be subjected to under Article 37 of the DSA [19].
It has been argued by Mökander et al. [62] that the EU regulation thereby sketches a de facto EU-wide ecosystem for auditing AI systems. However, Edwards [24] has argued that self-assessments before deployment and post-market monitoring – we go with Mökander et al. [62] to frame this as a form of internal auditing – do not provide sufficient oversight. Therefore, Edwards [24] has raised the question of the extent to which audits and assessments by third parties (regulator, research, and civil society) are necessary for sufficient oversight.
As this paper will demonstrate, access to elements such as the model and training data of the algorithmic system is crucial to facilitate external auditing. In this sense, we reiterate and strengthen, on the one hand, the argument of Casper et al. [16] that third-party black-box access — in which “auditors can only query the system and observe its outputs”, as opposed to “white-box access to the system’s inner workings” — is insufficient and, on the other, AlgorithmWatch’s concern that the fact that third-party access is lacking in the AIA is worrisome for the efficacy of the regulation [4]. Moreover, while the DSA includes third-party access by vetted researchers, scholars have argued that companies could leverage their market power against external auditors (i.e. audit capture) [50]. This is why we call for extended access for civil society and journalists to strengthen auditors positions and oversight by civil society during the implementation of both the DSA and for future regulations.
@@
In general, researchers wonder “who will audit internal auditors?” [22, 3, 24]. Therefore, we analyze whether the DSA and the AIA can create an effective, diverse EU audit ecosystem that is capable to protect the rights that they aim to protect. To do so, we examine the importance of third-party audits by researchers and civil society. We find that third-party audits are essential for the establishment of an AI auditing ecosystem in the EU that ensures accountability by compliance and oversight. We argue that the inclusion of such audits must be strengthened by considering the existing provisions of the DSA and the provisions of the AIA, which comes into force in 2024. The question of how audits conducted by the regulator should be included in an AI auditing ecosystem unfortunately goes beyond the scope of the paper, although researchers have justifiably insisted on the importance of this issue (see e.g. [40]).
@@
We contribute to the existing literature concerning third-party audits in three ways: (1) We define an AI audit ecosystem that accounts for compliance and oversight. (2) We demonstrate the existence of a regulatory gap as the DSA and AIA will not lead to a diverse AI audit ecosystem that includes civil society and, in the case of the AIA, even vetted researchers. (3) We emphasize that external audits by vetted, independent research centres, journalists and civil society organisations have to be part of that ecosystem. For this reason, we argue that the AI Act has lost an important opportunity to include should include broader access to AI systems, as it is provided in the case of the DSA, and that this gap should be addressed by future regulation and AIA amendments.
@@
First, we introduce the terms related to algorithmic audits and the different types of audits (section 2). We then analyze the current regulatory framework on third-party audits (section 3). The following section addresses third-party audit case studies and defines an AI audit ecosystem that ensures compliance and oversight. Furthermore, we show that third-party audits by researchers and civil society are vital for an AI audit ecosystem (section 4). Finally, we argue that the current EU legislation – the DSA and AIA in combination – will not be sufficient to create a diverse AI audit ecosystem that includes researchers and civil society representatives. We then make concrete proposals for future regulation, as well as for delegated acts and potential amendments for the DSA and the AIA (section 5).
@@ 2 Algorithm Audits: Terminology and Key Properties
In this section, our primary objective is to establish a robust foundation of terminology and key attributes inherent to algorithm audits. Consequently, we will distinguish audit types, examine the scope of audit types, and underline the role of sociotechnical system thinking in audits. These examinations are conducted in preparation for the analysis of the role of third parties in the DSA and AIA. Additionally, we will revisit these audit types in the discussion of creating a diverse AI audit ecosystem.
@@
The term “audit”  [^2] Sandvig et al. [82] have noted that “although the word ’audit’ may evoke financial accounting, the original audit studies were developed by government economists to detect racial discrimination in housing.” Later, they were introduced in the social science as field experiments to test for discriminatory behavior [33]. is used in various ways. For algorithm audits [^3] In the following, we will use the terms ”audit” and ”algorithm audit” interchangeably., some (e.g. Koshiyama et al. [48]) have provided a more technical definition of an algorithm audit, which is related to verification and compliance, and others (e.g. Sandvig et al. [82] and Metaxa et al. [59]) have used the term to describe a specific but still systematic targeted test of a particular aspect of a system (e.g., bias), originating from social science “audit studies.”
@@
Two examples demonstrate the various ways that the term has been used. Firstly, Bandy [8] identifies a research gap in audits of Twitter. Twitter is the most researched social media platform due to its formerly openly-accessible API [68]. This suggests a conceptual problem: What differentiates research on Twitter from an audit conducted on Twitter? In the second example, Metaxa et al. [59] defines audits as systematic external evaluations by third parties, serving as a form of activism with internal knowledge of the process or system being studied.
@@
We base our understanding of audits on Raji et al. [76] and Bandy [8] by defining “audit” as an empirical study that investigates algorithmic systems and evaluates performance relative to expected behavior as part of a broader accountability process. The focus of an audit is to examine for potential problematic behavior, which refers to system behavior with the potential to cause harm to individuals, groups, or society [8]. “Empirical studies” refers to qualitative and quantitative studies, as system expectations can be articulated and assessed in either mode [76]. We note that this definition is also agnostic to the time of the audit (before or after the deployment).
@@
In a systematic review of algorithm audits, Bandy [8] points out that racial discrimination remains a central focus. Thus, Bias audits address algorithmic discrimination based on race and other protected attributes, such as age, gender, socioeconomic status, religion, and intersectional identities. However, apart from representational and allocative harms related to discrimination, AI systems can pose other potential risks. Therefore, audits may need to have different scopes to address quality of service harms, interpersonal harms, and social system harms [85]. To address these harm types, audits such as robustness audits, interpretability and explainability audits, security audits, and privacy audits can be conducted [48].
@@
  • first-party audits: internal audits which are conducted within an organization that developed the AI system or product based on specific metrics, toolkits, and requirements.
  • second-party audits: conducted by contractors with the developer (audits-as-a-service), including consulting companies as the “big four,” as well as companies that specialize in audits. While maintaining a degree of independence from deployers, risk “audit washing” (see section 4.3.1) by potentially catering to their clients’ interests.
@@
First-party assessments and audits Second- and Third-party assessments and audits
Algorithmic risk assessment, algorithmic impact assessment Ethic-based auditing and AI safety auditing Internal conformity assessment and post-market monitoring Research and civil society algorithm audits, bias audits, adversarial audits Regulatory inspection, external conformity assessments and certification Algorithmic impact evaluations, sociotechnical audits, ecosystem audits
Example(s) Case study algorithmic impact assessment for data access in a healthcare context [37], data protection impact assessment [11] End-to-end framework for internal algorithmic auditing by Raji et al. [75] AIA conformity assessments, example procedure by [32]; Medical Device Regulation Application Procedure [18] Gender Shades [14], Audit of the Rotterdams’ welfare fraud ADM [20, 21], Audit Austrian AMS algorithm [5] UK Information Commissioner’s Office’s ‘Guidance on the AI auditing framework’[45] Stanford’s Impact evaluation of a predictive risk modeling tool [35], sociotechnical audit that assess police use of facial recognition[72]
Auditors Creators or commissioners of the algorithmic system, contractors such as consulting companies Creators or commissioners of the algorithmic system, contractors such as consulting companies Creators or commissioners of the algorithmic system, contractors such as consulting companies Researchers, investigative journalists, data scientists, NGOs, affected communities, or other stakeholders Regulator or other auditing and compliance professionals that have an institutional role Researchers, policymakers, investigative journalists, data scientists, NGOs, affected communities, or other stakeholders
Target Identification & Audit Scope Broad scope: focus on the risks and negative consequences of the deployment of an algorithmic system Narrow scope: focus predominantly on technical prevention and uncovering of specific harms (e.g. bias audit, safety audit) Broad scope: focus on an algorithmic system’s compliance with regulation and mostly non-technical and process-oriented Narrow scope: focus predominantly on technical prevention and uncovering of specific harms (e.g. bias audit) Broad scope: focus on an algorithmic system’s compliance with regulation Broad scope: assessing possible societal impacts of an algorithmic system on the users or population it affects after it is in use
@@
Auditor Independence Internal and thus no independence or contractual dependence Internal and thus no independence or contractual dependence Internal and thus no independence or contractual dependence External by independent organizations with no obligation External by independent institution, non-genuine independence when auditing administrative state-run ADMs External by independent organizations with no obligation
@@
Auditor Professionalization Formal: done by deployers, internal teams or accredited second-parties Formal: Done by deployers, internal teams or accredited second-parties Formal: done by deployers, internal teams or accredited second-parties Mostly informal: accreditation and cooperation with company possible , although basic access could be possible without accreditation (’right to scrape’) Formal: Accreditation Mostly informal: accreditation and cooperation with company possible (vetted researchers in DSA)
@@
When to audit Before deployment Before deployment Before deployment, post-market monitoring is possible After deployment After or before deployment, potentially an ongoing process After deployment
@@
Table 1: This table is inspired by and derived from [2, 59, 63, 76] and should give an overview of assessments and audits that we surveyed. The considerations of Raji et al. [76] were used to display the differences of these assessment and audit types. However, this table claims neither completeness nor complete adequacy, as this topic is a relatively emerging one with respect to algorithmic systems. At the same time, a systematic presentation of assessments and audits differences is not the focus of this paper, even though we conceptually argue and advocate a broad deployment of assessments and audits in the AI lifecycle within the AI audit ecosystem.
@@
Algorithms cannot be divorced from the contexts in which they are applied [97] and represent a complex network containing economic, environmental, political, and social interests and outcomes affecting their functioning. The entanglement of technical components and social elements leads to the typical acknowledgment of AI systems as sociotechnical systems [^4] Due to a broader understanding of their ecological impacts in recent years, [77] has added an “ecological” element to this notion, meaning that AI systems are socioecological-technical systems. Thus, it may be necessary to include ecological harms that AI causes in audits. [83] And, despite the common perception of audits as predominantly technical (see e.g. Koshiyama et al. [48]), internally conducted impact assessment as well as externally conducted impact evaluations and sociotechnical audits encompass a broader spectrum. They evaluate the impact of an algorithmic system on a population in a comprehensive framework that encompasses technical elements – most of the time limited technical elements – and include the social elements of the system [2]. We will come back to the assessment of sociotechnical systems in section 4.4.1.
@@
Ethics-based or AI safety internal auditing and adversarial audits by researchers and civil society have a more technical scope and tend to focus on specific harms, e.g. bias audits. Nevertheless, drawing a clear line of demarcation between narrow and broad scope of audit is difficult. While maintaining a however central focus on sociotechnical systems, it is essential to consider the data that substantiate claims concerning specific harms attributed to an AI system. An example of this difficult distinction is that Stahl et al. [88] includes bias mitigation assessments in the systematic review on impact assessments, which we have introduced as ethics-based audits or bias audits. Thus, these kind of audits could potentially be part of an impact assessment.
@@
Conformity assessments [62] and regulatory inspection [2] are systematic tests that evaluate if a system is compliant with a specification, standards, or a regulation. Regulatory inspections – conducted by an independent state-run body – are external to the deploying organization. Conformity assessments can be internal or external. In contrast to third-party audits, conformity assessments aim to achieve compliance, while third-party audits by researchers and civil society aim to prevent potential harm or uncover it by systematically evaluating the systems and checking the documents.
@@
We focus on third-party audits, especially those by researchers, academics, and journalists, as well as broader sociotechnical audits. [^5] These broader types of audits are also referred to as “ecosystem audits” by Birhane et al. [12]. Such audits are characterized by being conducted post-deployment through systematic evaluations – mostly a combination of different qualitative and quantitative methods – and reverse engineering of system behaviour, typically with little data and model access and independent of the organization deploying the algorithm [59, 22, 12]. Thus, they have no formal contractual relationship with the audit target [73].